OpenStudy (anonymous):

What percent of 60 is 12? Show your work??

OpenStudy (anonymous):

60% as a fraction is 3/5 If 3/5 is 12 5/5 is 12 x 5/3 =20 So the answer is 20 is this what you're asking? @britbrat4601

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

No problem :)

OpenStudy (lgbasallote):

divide 12 by 60 then multiply by 100% \[\implies \frac{12}{60} \times 100\text %\]